OpenStudy (anonymous):

By what approximate factor is the intensity of an earthquake with magnitude 5.4 greater than an earthquake with magnitude 5.3? 1.01 1.21 1.26 10.44

OpenStudy (anonymous):

5.4 = log(x/10^-4) 5.3 = log(x/10^-4)